OpenStudy (anonymous):

Suppose that the concentration of a bacteria sample is 30,000 bacteria per milliliter. If the concentration triples in 4 days, how long will it take for the concentration to reach 51,000 bacteria per milliliter? (In days)

OpenStudy (anonymous):

It's the 6th time I get the problem wrong. Each time I get a similar problem, I don't get it.

OpenStudy (anonymous):

@whpalmer4 @thomaster

OpenStudy (anonymous):

@e.mccormick @iambatman @jigglypuff314 Halp

OpenStudy (anonymous):

@AccessDenied @Ashleyisakitty @mathmale

jigglypuff314 (jigglypuff314):

@surjithayer @zepdrix @satellite73 @mathslover if you're not too busy could your guys try to help please? :3

OpenStudy (timaashorty):

I'm trying to work it out, although it's a bit confusing. Is there any answer choices?

OpenStudy (anonymous):

No there aren't. I know it's a Ca^x problem, just don't get what I'm doing wrong

OpenStudy (timaashorty):

Can you show what you've done so far?

OpenStudy (anonymous):

Well I tried setting it up. C is 30,000

Parth (parthkohli):

\[C(30,000)^4 = 120,000\]I think we can work around this way.

Parth (parthkohli):

Oh whoops, that 120,000 is supposed to be 90,000

OpenStudy (anonymous):

How did you set that up, m8?

Parth (parthkohli):

\(C\times a^x = \rm new~thing\) And \(a= 30,000\), \(x = 4\). It's given that the thing triples so the final quantity is \(90,000\)

Parth (parthkohli):

So we can solve for \(C\) from that and plug into\[C\times (30,000)^x = 51,000\]

Parth (parthkohli):

u alrite m8?

OpenStudy (anonymous):

Wouldn't C be 30,000? As the original thing? I thought A was the growth factor

Parth (parthkohli):

Oooh.

zepdrix (zepdrix):

Luce do you get unlimited guesses? or does it give you a new question after you've attempted this one? I have an answer but I'm not sure if I did it correctly :o

Parth (parthkohli):

OK, then use the equation\[30,000 \times a^4=90,000\]

OpenStudy (anonymous):

I get three or two tries I believe. After I get it wrong, I get a similar question. So unlimited tries at a type of problem, a couple at the specific problem.

Parth (parthkohli):

OK, I promise this one was right.

OpenStudy (anonymous):

\[a ^{4}=3\]

Parth (parthkohli):

Right... log em up.

OpenStudy (anonymous):

Natural log, amirite?

Parth (parthkohli):

any base would work.

zepdrix (zepdrix):

So I was attempting this,\[\Large\rm B(t)=30000(a)^{kt}\]Bacteria as a function of time. Triples every 4 days, so a=3, and then we just adjust the days.\[\Large\rm B(t)=30000(3)^{t/4}\]So that should give triple every 4 days. :o This is kinda the same thing Parth was doing I guess. His approach is more thorough though. And then just let \(\Large\rm B(t)=51000\) and solve for t.

OpenStudy (anonymous):

So wait, what I'm I setting this up as Parth? Kinda confused. Also trying Zep

Parth (parthkohli):

\[a^4 = 3\]\[\Rightarrow \log(a^4) = \log(3)\]\[\Rightarrow 4 \log(a) = \log(3)\]\[\Rightarrow \log(a) = \log(3)/4\]\[\Rightarrow a = 10^{\log(3)/4} = 10^{0.119}= 1.31\]All credits go to my calculator.

Parth (parthkohli):

Now that we have found \(a\), we just plug it into the original equation\[30,000\times (1.31)^x = 51,000\]

OpenStudy (anonymous):

Alright so, 1.31^x=1.7?

Parth (parthkohli):

Right.

OpenStudy (anonymous):

Log'ing them now

OpenStudy (anonymous):

(x)0.2700=0.5306

OpenStudy (anonymous):

So far so good?

Parth (parthkohli):

I'd think so, all right.

OpenStudy (anonymous):

x=1.9651

Parth (parthkohli):

All righty. So it'd take almost 2 days for 51,000.

OpenStudy (anonymous):

I did something wrong in the math. The answer was 1.93

Parth (parthkohli):

Hmm, seems like we lost significant digits in the process. Let's not calculate the value of any logarithm and take it down to the end. With that, we have the following:\[a^4 = 3\]\[\log (a) = \log(3)/4\]\[a = 10^{\log(3)/4}\]So plugging this thing into the original equation...\[\left(10^{\log(3)/4}\right)^x = 1.7\]Messy expression, yes, but we'll have to calculate with the exact value when dealing with exponents. I should have realized this earlier.

Parth (parthkohli):

Have you lost all of your attempts? If so, I'd like to apologize.

OpenStudy (anonymous):

Yes, don't worry it's okay. You want help me through the similar problem? My got it's the seventh time

Parth (parthkohli):

I hope I can. :)

OpenStudy (anonymous):

Suppose that the concentration of a bacteria sample is 20,000 bacteria per milliliter. If the concentration triples in 4 days, how long will it take for the concentration to reach 32,000 bacteria per milliliter? (In days)

OpenStudy (anonymous):

Is a^4=3 again?

Parth (parthkohli):

Oh, this question is very similar. Yes - that's right.

OpenStudy (anonymous):

I'm still confused as to where we went wrong in the math. The problem says to round all intermediate numbers to the nearest thousandth, is that it?

Parth (parthkohli):

Ah, that is the problem. We didn't know how to round-off. OK, we'll now do it to the nearest thousandth.

OpenStudy (anonymous):

So, 4log(a)=log(3)?

OpenStudy (anonymous):

Then log(a)=(log(3))/4

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

log(a)=0.119

Parth (parthkohli):

Exactly. (OK, not exactly - just rounded off :P)

OpenStudy (anonymous):

So how do I take the log out of log(a)?

Parth (parthkohli):

\[\rm \log(a) = something \Rightarrow a = 10^{\large something}\]Definition of log.

Parth (parthkohli):

That's how we did it in the previous question.

OpenStudy (anonymous):

so 10^0.119=1.315

Parth (parthkohli):

Seems good.

OpenStudy (anonymous):

a=1.315?

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

Then 20,000(1.315^x)=32,000

OpenStudy (anonymous):

1.315^x=1.6

Parth (parthkohli):

Right.

OpenStudy (anonymous):

(x)log(1.315)=log(1.6)

Parth (parthkohli):

Mhm, mhm.

OpenStudy (anonymous):

x=1.716, final answer being 1.71?

Parth (parthkohli):

I guess you should round-off the final answer to the nearest thousandth as well.

OpenStudy (anonymous):

Omg guess what

Parth (parthkohli):

I'm sorry I got another one wrong. :(

OpenStudy (anonymous):

No silly, you da man

Parth (parthkohli):

lulz k

OpenStudy (anonymous):

I think we did something slightly wrong in the math again though. I put 1.71 as the answer, though when you rounded off it should have been 1.72. lmao Anyways 1.71 was right

Parth (parthkohli):

Alrighty then!